How would I determine the mmoles of phosphoric acid used in an experiment?

experiment details:

phosphoric acid 85% in water
amount used: 1mL
molecular weight: ?
density: ?
Answer #1

Given that:-

Phosphoric acid: 85% in water (w/w)

(w/w means 85 g of phosphoric acid in 100 g of water)

Amount of solution used: 1ml

Molecular weight of phosphoric acid: 97.994 g/mol

Density of phosphoric acid: 1.88 g/ml

Answer:-

​​​​​Mass of phosphoric acid used to make solution is: 85g

Therefore volume of phosohoric acid used is:

Volume = mass/ density

= 85/1.88

= 45.21 ml

Also as the density of water is 1 g/ml therefore volume of water used is 100 only

Hence, total volume of solution is: 100 ml water+ 45.21 ml phosohoric acid = 145.21 ml

Now,

No. of moles of phosohoric acid in 145.21 ml of solution is given by:

No.ofmoles = givenmass/molarmass

= 85/97.994

= 0.867 moles

No. of moles of phosphoric acid in 1 ml of solution is:

= 0.867/145.21

= 0.00597 moles

Therefore no. of mmoles of phosohoric acid in 1 ml solution is:

= 0.00597/1000

= 0.00000597

= 5.97 * 10-6 mmoles

  

​​​​​​